# GLProtein: Global-and-Local Structure Aware Protein Representation Learning

### v1

@@ 4.2.3 Homology Detection, Fluorescence and Stability Prediction
Overview of homology detection. The task of predicting remote homology in proteins can be viewed as a classification problem at the molecular level. The objective is to input a protein sequence into the homology detection model, which then identifies the correct types of protein fold. In our paper, this presents a significant challenge with 1,195 distinct protein folds to classify. We utilize data sources from  hou2018deepsf and present the average accuracy achieved on the fold-level heldout set.
@@
Overview of fluorescence prediction. In the realm of protein science, fluorescence prediction is a vital task that involves estimating the fluorescence properties of proteins. This is a regression task where each input protein is mapped to a label measuring the most extreme circumstances in which the protein maintains its fold above a concentration threshold. We use the data from  rocklin2017global and use Spearman’s rank correlation coefficient as the metric.
@@
Overview of stability prediction. Stability prediction involves estimating the resilience of a protein’s structure under various environmental conditions, a critical factor in understanding its functional efficacy and therapeutic potential. This regression task focuses on predicting the intrinsic stability of proteins, which is essential for assessing their capacity to preserve their structural integrity under severe conditions. To assess the effectiveness of our model, we measure its performance using Spearman’s rank correlation coefficient across the entire test set rocklin2017global .
@@
4.2.4 Protein Function Prediction
Figure 6: Comparisons on semantic similarity inference. The best results are bolded, and the second-best results are underlined.
Methods MF BP CC Avg
@@ 4.2.4 Protein Function Prediction
Overview. Protein function prediction aims to assign biological or biochemical roles to proteins, and we also regard this task as a sequence classification task. Following KeAP zhou2023protein , we divide protein attributes into three groups: molecular function (MF), biological process (BP) and cellular component (CC), and report the Spearman’s rank correlation scores for each group.
@@
We also evaluate our GLProtein with 6 baseline models on protein-protein binding affinity estimation task, which is detailed in Appendix 7.2.
5 Failure Discussion
6 Conclusion and Future Work
In this work, we propose GLProtein, a general protein pre-training framework with global and local protein structure information. By integrating structure similarities between proteins, protein 3D distance encoding and amino acid molecule embeddings, GLProtien outperforms the previous protein representation learning model on most downstream applications, demonstrating the performance superiority of GLProtein. In the future, we aim to further enhance GLProtein’s capabilities by exploring novel avenues for incorporating multi-modal data sources, refining the model’s interpretability, and extending its applicability to a wider array of biological contexts.

### v2

@@ A.3 Homology Detection, Fluorescence and Stability Prediction
Overview of homology detection. The task of predicting remote homology in proteins can be viewed as a classification problem at the molecular level. The objective is to input a protein sequence into the homology detection model, which then identifies the correct types of protein fold. In our paper, this presents a significant challenge with 1,195 distinct protein folds to classify. We utilize data sources from  Hou et al. (2018) and present the average accuracy achieved on the fold-level heldout set.
@@
Overview of fluorescence prediction. In the realm of protein science, fluorescence prediction is a vital task that involves estimating the fluorescence properties of proteins. This is a regression task where each input protein is mapped to a label measuring the most extreme circumstances in which the protein maintains its fold above a concentration threshold. We use the data from  Rocklin et al. (2017) and use Spearman’s rank correlation coefficient as the metric.
@@
Overview of stability prediction. Stability prediction involves estimating the resilience of a protein’s structure under various environmental conditions, a critical factor in understanding its functional efficacy and therapeutic potential. This regression task focuses on predicting the intrinsic stability of proteins, which is essential for assessing their capacity to preserve their structural integrity under severe conditions. To assess the effectiveness of our model, we measure its performance using Spearman’s rank correlation coefficient across the entire test set Rocklin et al. (2017).
@@
A.4 Protein Function Prediction
Overview. Protein function prediction aims to assign biological or biochemical roles to proteins, and we also regard this task as a sequence classification task. Following KeAP Zhou et al. (2023a), we divide protein attributes into three groups: molecular function (MF), biological process (BP) and cellular component (CC), and report the Spearman’s rank correlation scores for each group.
Methods MF BP $C$ Avg
@@ A.4 Protein Function Prediction
Table 3: Comparisons on semantic similarity inference. The best results are bolded, and the second-best results are underlined.
@@
A.5 Protein-Protein Binding Affinity Estimation
Overview. In this task, we focus on assessing how well protein representations can predict changes in binding affinity caused by protein mutations. This regression task involves associating each protein pair with a numerical value. Following the methodology described in  Unsal et al. (2022), we employ Bayesian ridge regression on the outcomes of element-wise multiplication of representations derived from pre-trained protein models. This approach is designed to enhance the accuracy of binding affinity predictions. We used the SKEMPI dataset from  Moal and Fernández-Recio (2012) and reported the mean square error of 10-fold cross-validation.
Baselines. We evaluate our model compared with six baselines. Specifically, we employed PIPR Chen et al. (2019), ProtBert Elnaggar et al. (2021), ESM-2 Rives et al. (2021); Verkuil et al. (2022); Hie et al. (2022), SaProt Su et al. (2023), OntoProtein Zhang et al. and KeAP Zhou et al. (2023a). PIPR is a siamese-residual-RCNN-based model for multifaceted protein–protein interaction prediction. ProtBert is a BERT-like model pre-trained on UniRef100 Suzek et al. (2007, 2015). ESM-2 feature a transformer architecture pre-trained on the representative sequences from UniRef50 Suzek et al. (2007, 2015). SaProt is the most recent structure-based protein language model. OntoProtein and KeAP are the most recent knowledge-based pre-training methodologies.
Methods Affinity($\downarrow$)
PIPR 0.63
ProtBert 0.58
ESM-2 0.48
SaProt 0.58
OntoProtien 0.59
KeAP 0.52
GLProtein 0.52
Table 4: Comparisons on protein-protein binding affinity prediction, with the best result bolded and the second best underlined. The notion $\downarrow$ signifies a preference for lower values, reflecting a superior predictive performance in this context.
Parameters Resouces Pre-training Inference
(40 examples)
ProtBert 400M A single TPU Pod V3-512 400k steps 2.02s
OntoProtein 400M 4 NVIDIA 48G A6000 GPUs 3 Days (continue pertaining on ProtBert) 1.91s
KeAP 400M 4 NVIDIA 48G A6000 GPUs 3 Days (continue pertaining on ProtBert) 1.94s
SaProt 650M 64 NVIDIA 80G A100 GPUs 3 Months 3.02s
ESM-2 650M - - 2.45s
GLProtein 400M 4 NVIDIA 48G A6000 GPUs 3 Days (continue pertaining on ProtBert) 1.93s
Table 5: Comparison of the number of parameters, resources, pre-training time, and inference time for GLProtein and baselines.
Results. Table 4 compares several methods of predicting the binding affinity of protein interactions, where a lower score indicates superior performance. GLProtein outperforms PIPR, ProtBert, SaProt and OntoProtein. It also shows the competitive performance of KeAP and ESM-2.
A.6 Ablation Study
Figure 6: Left: Ablations of three proposed approaches. Long-range P@L/2 results are reported for contact prediction. Right: Ablations of three proposed approaches. F1 scores are reported for protein-protein interaction tasks.
We investigated the effects of employing diverse protein structure information fusion strategies. First of all, the exclusion of the global structure information modelling component (representation as "w/o triplet" in Figure 6) resulted in varying degrees of performance deterioration across contact prediction and protein-protein interaction prediction tasks. This observation suggests that our global structure similarities through protein triplet contrastive learning stand out as a more efficacious choice. Subsequently, upon removing the proposed substructure-based molecular encoding from the local protein structure information component (denoted as "w/o aa" in Figure 6), we noted a decline in performance by approximately 2.5% and 8% for contact prediction and protein-protein interaction tasks, respectively. This underscores the essential role of substructure-based molecular encoding within our proposed methodologies. Finally when the protein 3D distance encoding was omitted from the local structure information modelling component (indicated as "w/o coord" in Figure 6), a similar trend of performance degradation was observed, further emphasizing the indispensability of this strategy within our architectural framework.
A.7 Parameter Sensitivity Study
In this section, we explore the impact of the parameters in the model on the final performance of our protein model. We experimented with the number of protein samples in the protein local structure information modelling component and the coefficient of contrastive learning loss for the protein triplet, respectively.
As shown in Figure 7, we test the number of protein samples from 1 to 4 on the contact prediction task. We observe that as the number of protein samples increases, the performance of our model improves to varying degrees in short-, medium- and long-range contact prediction. This also shows that our proposed protein triplet approach indeed enables the protein language model to capture the structural similarity features among proteins. Due to computational and memory cost considerations, we ended up constructing 4 protein positive samples and 4 protein negative samples for each protein.
As shown in Figure 8, we test the value of the coefficient $\alpha$ of contrastive learning loss for the protein triplet. We divided the experiment into 6 groups and set the values of $\alpha$ to 0.1, 0.3, 0.5, 1, 3, and 5. Then, we evaluated them using the protein-protein interaction prediction task on SHS27k, SHS148k, and STRING datasets, respectively. We observe that the model achieves the best performance when the value of $\alpha$ is set to 1. Thus, we choose $\alpha=1$ in this paper as our model’s setting.
Structure Evolutionary Engineering
$S$-Q3 $S$-Q8 Contact Homology Fluorescence Stability
SaProt 0.51 0.45 0.37 0.12 0.25 0.46
LSTM 0.75 0.59 0.26 0.26 0.67 0.69
Transformer 0.73 0.59 0.25 0.21 0.68 0.73
ResNet 0.75 0.58 0.25 0.17 0.21 0.73
ESM-2 0.70 0.54 0.43 0.10 0.30 0.65
LM-GVP 0.69 0.50 0.43 0.20 0.64 0.69
GearNet 0.71 0.55 0.44 0.25 0.67 0.78
ProtBert 0.82 0.67 0.35 0.29 0.61 0.73
OntoProtein 0.82 0.67 0.40 0.24 0.65 0.74
KeAP 0.82 0.67 0.45 0.29 0.67 0.75
GLProtein 0.82 0.68 0.48 0.28 0.67 0.81
Table 6: Results on TAPE Benchmark. $S$ is a secondary structure task that is evaluated in CB315. In contact prediction, we test medium- and long-range using P@L/2 metrics. In protein engineering tasks, we test fluorescence and stability prediction using Spearman’s $\rho$ metric.
Figure 7: Parameter sensitivity study on the number of protein samples in the local structure information component.
Figure 8: Parameter sensitivity study on the value of the coefficient $\alpha$ of contrastive learning loss for the protein triplet in the local structure information component.
Task Epoch Batch Size Warmup Ratio Learning Rate Freeze Bert Optimizer
Contact 5 8 0.08 3e-5 False AdamW
Homology 10 32 0.08 4e-5 False AdamW
Stability 5 64 0.08 1e-5 False AdamW
$S$-Q3 5 32 0.08 3e-5 False AdamW
$S$-Q8 5 32 0.08 3e-5 False AdamW
Fluorescence 15 64 0.10 1e-3 True Adam
Table 7: Hyper-parameters for fine-tuning.
Our Proposed GLProtein
ESM-2
KeAP
ProtBert
Figure 9: Embedding visualizations of GLProtein, ESM-2, KeAP and ProtBert on SCOPe database.
A.8 Visualization of Protein Representations
To facilitate a more intuitive comparison, we utilize t-SNE to visualize the protein representations produced by GLProtein, ESM2, KeAP, and ProtBert. The visualization results, based on the non-redundant subset ($PIDE\leq 40\%$) of the SCOPe database Chandonia et al. (2019), are illustrated in Figure 9. As depicted in this figure, the representations for alpha and beta proteins generated by GLProtein are distinctly separated, whereas those produced by ESM-2, KeAP, and ProtBert are more closely intertwined.
A.9 Time Complexity Analysis
We provide a more specific complexity analysis as follows: protein encoder operates at approximately $O(L^{2}d)$, where $L$ is the length of protein sequence and $d$ is the embedding dimension. Triplet protein sampling operates at approximately $O(L^{3})$, reducing the complexity to $O(L^{2})$ by TM-Vec. Triplet loss operates at approximately $O(3Ld)\rightarrow O(Ld)$. Protein 3D distance encoding operated at approximately $O(KL^{2}d)$, where $K$ is the number of Gaussian Basis kernels. Substructure-based molecular encoding operates at approximately $O(Ld)$. Protein decoder operates at approximately $O(L^{2}d)$. Total computation cost operated at $O_{total}=O_{encoder}+O_{global}+O_{local}+O_{decoder}=O(L^{2}d)+O(L^{2})+O(Ld)+O(KL^{2}d)+O(L^{2}d)=O((K+1)L^{2}d)$.
A.10 Hyper-parameters for Fine-tuning
The hyper-parameters for fine-tuning are provided in the Table 7. Specifically, we follow the hyper-parameter settings in GNN-PPI Lv et al. (2021) for PPI prediction. For protein binding affinity prediction and semantic similarity inference, we follow the fine-tuning configuration in PROBE Unsal et al. (2022).